• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name olivier

    Log entry time 17:20:52 on October14,2002

    Entry number 88175

    keyword=41 uA !

    The laser was changed, and we now get 41 uA. At this current, the raster was increased to 2x2 mm2 on the dummy target. We will go back to 1x1 mm2 when we switch to LD2 target.